A particle is projected vertically upward with an initial velocity of 60m/sec find the distance and displacement by the particle in 
A) 6sec. 
B) 4sec
C) 8sec

Grade:11

2 Answers

Vikas TU
4 years ago
Hiii
Initial velocity = 60 m/s
Final velocity = 0m/sec
v = u+at 
So t = 6 sec 
S = ut + 1/2 at 2
Find the distance travelled by this formula by putting value of different t and acceleration due to gravity = - 10 m/sec 2
S = ut+ 1/2 at2
 
Khimraj
4 years ago
Time at which it will reach its maximum is initial velocity/g = 6 sec 
distance and displacement by the particle in 
A) 6sec. 

Distance = Displacement = Maximum height = ut-{1/2}gt^2 = 180m